Yegnaw Company produces and sells two types of yoga-training products: how-to videotapes and a basic equipment set (blocks, strap, and small pillows).Last year, Yegnaw sold 14,500 videos and 7,250 equipment sets. Information the two products is as follows;


                                              Video set                                                Equipment

 

Price                                Br.17.40                                                     Br.21.75

Variable costs per unit       5.80                                                              8.70

Total fixed costs are Br.101, 500

 Required: Answer the following

1. What is the sales mix of videos and equipment sets

2. Compute weighted average contribution margin

 3. Compute the break-even quantity of each product.

 4. What is weighted average contribution margin ratio

5. What is the overall break-even sales revenue

 

 

 

 

 

 


1
Expert's answer
2022-03-21T18:18:03-0400
  1. 14500+7250 = =21750
  2. (101500 - 17.50) = 101482.50
  3. 14500/3= 4833.33
  4. (101500 - 17.50) = 101482.50
  5. 14500+7250 = 21750